OpenStudy (anonymous):

FAN + MEDAL Which ordered pairs are solutions to the equation? y = 2x + 5 A. (15, 5) and (2, 9) B. (5, 15) and (2, 9) C. (5, 10) and (2, 9) D. (5, 15) and (9, 2)

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

What is 2 times 5 plus 5?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

\[2 \times 5 + 5 = 15\] That makes the answer either B or D.

OpenStudy (anonymous):

What is \[2 \times 2 + 5?\]